# ACPO: AI-Enabled Compiler-Driven Program Optimization

### 15 Dec 2023 v1

@@ 5.3. Combined Results
Finally, we deploy both ACPO models in combined mode and experimentally observe the benefits on Polybench and Cbench. The combined results have both ACPO’s optimized loop unrolling and function inlining when used in LLVM’s O3 pass pipeline. An inspection for O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling (Ashouri et al., 2022).
Table 7. ACPO Combined Function Inlining & Loop Unrolling Performance (Polybench)
Benchmark Time O3 Var O3 Size O3 Time ACPO Var ACPO Size_ACPO Size Bloat Speedup
2mm 3.472 0.120 72608 3.530 0.023 138144 1.903 0.984
3mm 5.186 0.426 72608 5.311 0.309 203680 2.805 0.976
atax 0.738 0.101 72720 0.748 0.034 72696 1.000 0.987
bicg 1.374 0.081 72664 1.345 0.254 72664 1.000 1.022
cholesky 2.478 0.022 72752 2.330 0.027 72776 1.000 1.063
correlation 2.625 5.276 72728 2.619 0.539 138240 1.901 1.002
covariance 2.743 0.085 72648 2.719 0.988 138184 1.902 1.009
deriche 0.273 0.788 72648 0.245 0.088 138152 1.902 1.115
doitgen 0.897 0.007 72680 0.886 0.015 138192 1.901 1.013
durbin 1.359 0.244 72696 1.395 0.270 72696 1.000 0.974
floyd-warshall 28.314 0.034 72624 18.946 0.011 138160 1.902 1.494
gemver 2.702 0.154 72616 2.772 0.213 72616 1.000 0.975
gesummv 1.503 0.039 72616 1.511 0.019 72616 1.000 0.995
gramschmidt 3.990 0.545 72752 3.988 0.787 138256 1.900 1.000
lu 5.184 0.427 72696 5.207 0.914 72720 1.000 0.996
ludcmp 5.083 0.255 72720 4.767 0.283 72720 1.000 1.066
mvt 2.053 0.042 72608 2.027 0.040 72608 1.000 1.013
nussinov 3.840 0.089 72696 3.270 0.067 72696 1.000 1.174
symm 2.728 0.046 72648 2.544 0.676 138184 1.902 1.072
syr2k 3.136 0.538 72608 3.171 0.971 138144 1.903 0.989
syrk 1.504 0.382 72608 1.452 0.765 138144 1.903 1.036
trisolv 1.149 0.028 72648 0.933 0.098 72648 1.000 1.232
trmm 1.374 0.496 72648 1.409 0.288 72648 1.000 0.975
Geomean - - - - - - 1.383 1.045
Table 8. ACPO Combined Function Inlining & Loop Unrolling Performance (Cbench)
Name ACPO-FI ACPO-LU ACPO Combined
Size Bloat Speedup Size Bloat Speedup Size Bloat Speedup
automotive_bitcount 1.000 0.999 1.001 1.003 1 1.001
automotive_qsort1 1.000 0.994 1.000 0.982 1.91 0.989
automotive_susan_c 1.000 0.999 3.685 1.035 19.797 1.066
automotive_susan_e 1.000 1.023 3.685 1.019 19.797 1.026
automotive_susan_s 1.000 0.980 3.685 1.038 19.797 0.965
bzip2d 1.441 0.991 2.224 1.010 3.541 1.014
bzip2e 1.441 1.007 2.350 1.001 3.546 1.015
consumer_jpeg_c 1.288 1.017 5.385 0.986 5.086 1.027
consumer_jpeg_d 1.289 0.923 5.400 1.073 4.807 1.026
consumer_lame 1.437 1.027 4.501 1.030 4.729 1.034
consumer_mad 1.217 1.071 4.321 0.990 3.193 0.968
consumer_tiff2bw 1.180 1.002 4.502 0.943 2.471 0.987
consumer_tiff2rgba 1.180 1.250 4.681 1.036 2.657 1.012
consumer_tiffdither 1.180 1.046 1.213 0.967 2.655 1.002
consumer_tiffmedian 1.180 1.017 1.000 0.970 3.021 0.943
network_dijkstra 1.000 0.973 1.000 0.990 1 1.037
network_patricia 1.000 0.970 1.000 0.974 1 1.004
office_ghostscript 1.275 1.000 1.000 1.000 1.12 1.001
office_ispell 0.996 1.000 1.000 1.000 1 1
office_rsynth 1.000 1.000 2.024 1.010 2.024 1.011
office_stringsearch1 1.000 1.006 1.000 1.000 1 1
security_blowfish_d 1.000 1.002 1.000 0.978 1 0.965
security_blowfish_e 1.000 1.003 1.000 1.015 1 0.981
security_pgp_d 1.861 0.993 1.000 0.991 13.144 1
security_pgp_e 1.861 1.017 1.000 1.006 13.144 1.006
security_rijndael_d 1.000 0.990 3.823 0.966 1 0.966
security_rijndael_e 1.000 1.001 3.823 0.974 1 0.995
security_sha 0.999 1.270 1.000 1.174 0.999 1.501
telecom_adpcm_c 1.000 1.000 1.000 1.008 1 1.038
telecom_adpcm_d 1.000 0.997 1.000 1.132 1 1.138
telecom_CRC32 1.000 1.094 1.000 1.042 1 1.112
telecom_gsm 1.810 1.076 1.815 1.15 1.81 1.117
Geomean 1.152 1.021 1.818 1.01 2.442 1.024
The results are shown in Tables 7 and 8. We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively. LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Module and Function level for Inline and Loop level for Loop Unroll, is outside the scope of this work. Instead, we showcase the overall benefits of each benchmark. For instance, Table 8 showcases the performance gains using ACPO-FI, ACPO-LU, and the combined version with separate columns.
@@ 41  Loop Size = 73
One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (show on Table 8) using ACPO’s Inliner, now with 50% performance gain. There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 5. Additionally, each corresponding line at Listing 6 shows the decisions made by LLVM’s O3. Note that, partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 6. To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 5. However, evaluations showed that still ACPO was outperforming O3 by a margin of 7% on unrolling this particular benchmark. Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time. Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism as it is less than or equal to the calculated cost. This example showcases the superiority of using 30/13 well-representative loop/function features we designed in training ACPOLUModels to deduce the kind of complex correlation between a code segment and the value of unroll count and the Boolean decision of inlining a callsite to optimize the performance.
@@ 6. Discussion
ACPO is the first scalable framework for bringing ML into compilers. It provides the ability to reproduce models on different architecture very fast and the methodology lays out the steps to regenerate data, retrain a new model, and easily deploy with LLVM. Compiler autotuning and optimization space have benefited from ML for almost two decades. As we pointed out earlier, there are many recent works leveraging ML, however, the goal of ACPO is to provide a novel flexible framework for the LLVM community to leverage ML on any optimization passes of interest with minimal efforts. Nevertheless, there are a number of challenges and shortcomings we would like to mention here.
Integration and dependencies
ACPO attempts to alleviate the challenging task of supply chain management and the use of licensed tools by separating the ML framework from the compiler. We showed both the integration of PyTorch and TensorFlow among other moving objects, i.e., Autotuner, into our framework, however, upstreaming would remain challenging and requires community efforts to shape the next generations of ML-enabled compilers. We also provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers. As mentioned in Section 3.6, ACPO uses Python named pipes to perform interprocess communication. Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table 3). Developers can switch between these modes when building the compiler. The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
Multi-objective optimization
ACPO optimizes the performance of a code region with intelligent unrolling decisions. However, in a number of cases, we experimentally observe that the emitted code size is also increased. This is a(n) (in)direct result of Inliner and Loop Unroll which increase the number of instructions and thus, the code size. Although this is outside the scope of the current work, an enhanced version may also benefit from taking into account both of the objectives in its exploration strategy by identifying the Pareto optimality (Hoste and Eeckhout, 2008; Ashouri et al., 2013; Moreau et al., 2018; Silva et al., 2021).
@@ 7. Conclusion
In this paper, we presented ACPO, the first scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM. We demonstrated that using ACPO framework, we are able to enable Loop Unroll pass to benefit from a seamless integration of an ML model, i.e., ACPOLUModel, for which we experimentally achieve on average a 4%, 3%, 5.4%, and 0.2% performance speedup over LLVM’s O3 on Polybench, Coral-2, CoreMark, and Graph-500, respectively. Additionally, we showcased a scenario where we leverage an existing ML model to optimize Function Inlining pass and similar to the original model’s performance, we experimentally obtained 2% speedup wrt LLVM’s O3 when deployed on CBench and we showed that by combining both models, the benefits increase up to 4.5% and 2.4% on Polybench and Cbench, respectively. Future works will be focused on extending ACPO framework by enabling a number of other optimization passes inside LLVM’s pipeline and leveraging innovative software characterization methods to be employed.
Appendix A Appendix —  ACPO Headers
1class ACPOModel {
2public:
3  ACPOModel(...);
4  ~ACPOModel();
5  void setMLIF(std::shared_ptr<ACPOMLInterface> ML);
6  std::shared_ptr<ACPOMLInterface> getMLIF();
7  void addRequiredResultField(std::string Name, Type::TypeID &ID);
8  std::unique_ptr<ACPOAdvice> getAdvice();
9  void addFeature(int64_t ID, Constant *Val);
10  void setCustomFeatures(const std::vector<std::pair<std::string, Constant *>> &FeatureValues);
11
12protected:
13  virtual void prepareModelInput();
14  virtual bool runModel(std::unique_ptr<ACPOAdvice> &Result);
15};
Listing 7: ACPOModel Header
1class ACPOMLInterface {
2public:
3  ACPOMLInterface();
4  virtual ~ACPOMLInterface();
5  virtual bool loadModel(std::string ModelSpecFile);
6  virtual bool freeModel(std::string ModelName);
7  virtual bool registerModel(std::string ModelName, int NumFeatures, int NumOutputs);
8  virtual bool registerFeature(std::string ModelName, std::string FeatureName, int Index);
9  virtual bool registerOutput(std::string ModelName, std::string OutputName, std::string OutputType);
10  virtual bool setCustomFeature(std::string ModelName, uint64_t FeatureID, int FeatureValue);
11  ...
12  virtual bool setCustomFeature(std::string ModelName, std::string FeatureName, bool FeatureValue);
13  virtual bool setCustomFeatures(std::string ModelName, const std::vector<std::pair<uint64_t, std::string>> &FeatureValues);
14  virtual bool setCustomFeatures(std::string ModelName, const std::vector<std::pair<std::string, std::string>> &FeatureValues);
15  ...
16  virtual bool runModel(std::string ModelName);
17  virtual std::string getOutputType(std::string ModelName, std::string OutputName);
18  virtual int getModelResultI(std::string OutputName);
19  ...
20  virtual bool getModelResultB(std::string OutputName);
21  ...
22  virtual int getStatus();
23  virtual bool closeMLInterface();
24};
Listing 8: ACPOMLInterface Header

### 11 Mar 2024 v2

@@ 5.2. Combined Results
Finally, we deploy both ACPO models in combined mode and experimentally observe the benefits on Polybench and Cbench. An inspection for O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling (Ashouri et al., 2022).
Table 2. ACPO Experimental Results (Polybench)
Benchmark ACPO-LU ACPO Combined
Size Bloat % Speedup Size Bloat % Speedup
2mm 1.903 0.984 1.903 0.984
3mm 1.904 0.983 2.805 0.976
adi 1.0002 1.012 1.38 1.04
atax 0.999 0.985 1 0.987
bicg 1 1.02 1 1.022
cholesky 1.001 1.102 1 1.063
correlation 1.005 0.957 1.901 1.002
covariance 1.001 1.003 1.902 1.009
deriche 1.0018 1.078 1.902 1.115
doitgen 1.0001 1.011 1.901 1.013
floyd-Warshall 1.0004 1.493 1 0.974
gemver 1.0004 0.958 1.902 1.494
gesummv 1 1.143 1 0.975
gramschmidt 1 0.91 1 0.995
lu 1.001 0.978 1.9 1
ludcmp 1.0008 1.047 1 0.996
mvt 1 1.046 1 1.066
nussinov 2.803 1.171 1 1.013
seidel-2d 1.0005 1.0008 1 1.174
symm 1.0004 1.06 1.902 1.072
syr2k 1.0009 0.991 1.903 0.989
syrk 1.0005 1.002 1.903 1.036
trisolv 1 1.223 1 1.232
trmm 1.0004 0.946 1 0.975
Geomean 1.10193 1.040725 1.383 1.045
@@ 41  Loop Size = 73
Table 3. ACPO Experimental Result (Cbench)
Sz%: Size bloat, Sp: Speedup. Multicolumns 1, 3, and 4 are wrt. O3 and 2, is ACPO-FI compared against MLGO (Trofin et al., 2021).
Benchmark ACPO-FI wrt. MLGO ACPO-LU Combined
SzSp SzSp SzSp SzSp
auto_bitcount 1 0.999 1 1.005 1.001 1.003 1 1.001
auto_qsort1 1 0.994 1 1.003 1 0.982 1.91 0.989
auto_susan_c 1 0.999 1.567 1.122 3.685 1.035 19.797 1.066
auto_susan_e 1 1.023 1.567 1.04 3.685 1.019 19.797 1.026
auto_susan_s 1 0.98 1.567 1.008 3.685 1.038 19.797 0.995
bzip2d 1.441 0.991 1.239 1.021 2.224 1.01 3.541 1.014
bzip2e 1.441 1.007 1.189 1.041 2.35 1.001 3.546 1.015
cons_jpeg_c 1.288 1.017 1.2 1.005 5.385 0.986 5.086 1.027
cons_jpeg_d 1.289 0.923 1.19 1.335 5.4 1.073 4.807 1.026
cons_lame 1.437 1.027 1.137 1.007 4.501 1.03 4.729 1.034
cons_mad 1.217 1.071 1.16 1 4.321 0.99 3.193 0.968
cons_tiff2bw 1.18 1.002 1.351 1.23 4.502 0.943 2.471 0.987
cons_tiff2rgba 1.18 1.25 1.356 1.11 4.681 1.036 2.657 1.012
cons_tiffdither 1.18 1.046 1.14 1.014 1.213 0.967 2.655 1.002
cons_tiffmedian 1.18 1.017 1.048 1.056 1 0.97 3.021 0.983
network_dijkstra 1 0.993 1 0.996 1 0.99 1 1.037
network_patricia 1 0.98 1.052 0.99 1 0.974 1 1.004
office_ghostscript 1.275 1 1.16 1.04 1 1 1.12 1.001
office_ispell 0.996 1 1.17 1.02 1 1 1 1
office_rsynth 1 1 1.12 1.04 2.024 1.01 2.024 1.011
office_strings 1 1.006 1 1 1 1 1 1
secur_blowfish_d 1 1.002 1 1.002 1 0.978 1 0.965
secur_blowfish_e 1 1.003 1.39 1.029 1 1.015 1 0.981
security_pgp_d 1.861 0.993 1.12 1 1 0.991 13.144 1
security_pgp_e 1.861 1.017 1.12 1.042 1 1.006 13.144 1.006
secur_rijndael_d 1 0.99 1.39 1.039 3.823 0.966 1 0.966
secur_rijndael_e 1 1.001 1.09 1.035 3.823 0.974 1 0.995
security_sha 0.999 1.27 1.009 1.132 1 1.174 0.999 1.501
tele_adpcm_c 1 1 0.996 1.001 1 1.008 1 1.038
tele_adpcm_d 1 0.997 1 1 1 1.132 1 1.138
telecom_CRC32 1 1.094 1 1.066 1 1.042 1 1.112
telecom_gsm 1.81 1.076 1.05 1.04 1.815 1.15 1.81 1.117
Geomean 1.152 1.021 1.161 1.041 1.818 1.01 2.442 1.024
The combined results are shown in Tables 2 and 3. We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively. Note that we don’t deploy ACPO-FI on Polybench as it consists of small nested-loop kernels which might not directly benefit from an optimized inliner, however, the experimental result reveals that indirectly it can. The underlying reason is that by it can provide more opportunities for the subsequent passes registered under O3 (Ashouri et al., 2022). LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Call Graph vs Function level.
One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (show on Table 3) using ACPO’s Inliner, now with 50% performance gain. There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 5. Additionally, each corresponding line at Listing 6 shows the decisions made by LLVM’s O3. Note that, partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 6. To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 5. However, evaluations showed that still ACPO was outperforming O3 by a margin of 7% on unrolling this particular benchmark. Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time. Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism.
@@ 6. Discussion
ACPO provides a class of features at varying LLVM scopes, therefore, new passes can utilize the already available features with minimal or no efforts. It also provides the flexibility of supporting other types of features, i.e., Graph-based, embeddings, etc. Additionally, we provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers. As mentioned in Section 3.5, ACPO uses Python named pipes to perform interprocess communication. Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table 1). Developers can switch between these modes when building the compiler. The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
@@ 7. Conclusion
In this paper, we presented ACPO, the first scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM. We demonstrated that using ACPO framework, we are able to enable Loop Unroll and Function Inline passes to benefit from a seamless integration of ML models, i.e., ACPOLUModel and ACPOFIModel, for which we experimentally achieved on average 4% and %4.5 on Polybench, and 2.1% and 2.5% on Cbench. ACPO provides a generalized feature collection class for which compiler developers can add to their instantiated ACPOModel class of their need and provides utility classes to use other types of features as well. Future works will be focused on extending ACPO framework by enabling other features and optimization passes inside LLVM’s pipeline and providing innovative software characterization methods using large Languages Models (LLMs).